# Partial Label Learning with a Partner

### v1

@@ Introduction
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News
PL-AGGD 0.063 $\pm$ 0.010 0.690 $\pm$ 0.020 0.451 $\pm$ 0.023 0.610 $\pm$ 0.012 0.521 $\pm$ 0.004 0.605 $\pm$ 0.002
@@
0.073 $\pm$ 0.011 $\bullet$ 0.698 $\pm$ 0.023 $\bullet$ 0.380 $\pm$ 0.013 $\bullet$ 0.542 $\pm$ 0.013 $\bullet$ 0.492 $\pm$ 0.003 $\bullet$ 0.463 $\pm$ 0.002 $\bullet$
@@
0.073 $\pm$ 0.006 $\bullet$ 0.721 $\pm$ 0.024 $\circ$ 0.471 $\pm$ 0.016 $\bullet$ 0.664 $\pm$ 0.012 $\bullet$ 0.521 $\pm$ 0.004 $\bullet$ 0.608 $\pm$ 0.003 $\bullet$
$O$ 0.071 $\pm$ 0.001 $\bullet$ 0.721 $\pm$ 0.004 $\circ$ 0.470 $\pm$ 0,020 $\bullet$ 0.663 $\pm$ 0.011 $\bullet$ 0.522 $\pm$ 0.003 $\bullet$ 0.605 $\pm$ 0.002 $\bullet$
$P$ 0.076 $\pm$ 0.010 0.717 $\pm$ 0.020 0.473 $\pm$ 0.017 0.668 $\pm$ 0.014 0.534 $\pm$ 0.005 0.609 $\pm$ 0.002
Table 4: Ablation study of PLCP coupled with PL-AGGD. $\bullet$/$\circ$ indicates whether PL-AGGD-PLCP is statistically superior/inferior to its degenerated version according to pairwise $t$-test at significance level of 0.05.
Comparison with Stand-alone Methods
Experimental Settings
@@
We conduct experiments on six real-world partial label data sets collected from several domains and tasks, including FG-NET (Panis et al. 2016) for facial age estimation, Lost (Cour et al. 2009), Soccer Player (Zeng et al. 2013) and Yahoo!News (Guillaumin, Verbeek, and Schmid 2010) for automatic face naming, MSRCv2 (Liu and Dietterich 2012) for object classification and Mirflickr (Huiskes and Lew 2008) for web image classification. The details of the data sets are summarized in the Supplementary.
For facial age estimation task, the ages annotated by crowd-sourced labelers are considered as each human face’s candidate labels. For automatic face naming task, each face scratched from a video or an image is presented as a sample while the names extracted from the corresponding titles or captions are its candidate labels. For object classification task, image segmentations are taken as instances with objects appearing in the same image as candidate labels.
Performance on Real-World Data Set
It is noteworthy that the number of average label of FG-NET is quite large, which could cause quite low classification accuracy for all approaches. The common strategy is to evaluate the mean absolute error (MAE) between the predicted age and the ground-truth one. Specifically, we add another two sets of comparisons on FG-NET w.r.t. MAE3/MAE5, which means that test samples can be considered to be correctly classified if the difference between the predicted age and true age is no more than 3/5 years. The results of these two comparisons are shown in the Supplementary due to the page limit. Table 6 summaries the classification accuracy with standard deviation of each approach on real-world data sets, where we can observe that
@@
  • State-of-the-art (SOTA) approaches, such as PL-CL, PL-AGGD, LALO and SURE, can also be significantly improved by PLCP on all data sets. For instance, on FG-NET the performance of SURE can be improved by 45%, and on FG-NET(MAE3) the performance of LALO can be improved by 5%. Additionally, PL-AGGD can be improved by 5.10% and PL-CL can be improved by 3.61%. The partner classifier’s non-candidate label information effectively and significantly aids disambiguation, leading to outstanding performance of the PLCP framework.
@@
Comparison with Deep-learning Based Methods
Experimental Settings
We conduct experiments on two benchmarks CIFAR-10 and CIFAR-100 (Krizhevsky, Hinton et al. 2009), and following the settings in (Wang et al. 2022; Lv et al. 2020), we generate false positive labels by flipping negative labels $\hat{y}\neq y$ for each sample with a probability $q=P(\hat{y}\in\mathcal{Y}|\hat{y}\neq y)$, where $y$ is the ground-truth label. In other words, the probability of a false positive is uniform across all $l-1$ negative labels. We combine the flipped ones with the ground-truth label to create the set of candidate labels. Specifically, $q$ is set to 0.1, 0.3 and 0.5 for CIFAR-10 and 0.01, 0.05 and 0.1 for CIFAR-100. Five independent runs are performed with the the average accuracy and standard deviation recorded (with different seeds).
Performance on CIFAR-10 and CIFAR-100
Table 2 presents the classification accuracy with standard deviation of each approach on CIFAR-10 and CIFAR-100, where we can find that
@@
Further Analysis
Improvement of the Disambiguation Ability
Transductive accuracy (i.e., classification accuracy on training samples) reflects the disambiguation ability of a PLL approach (Wang, Zhang, and Li 2022; Cour, Sapp, and Taskar 2011; Zhang, Zhou, and Liu 2016). In order to validate whether PLCP can correct some mislabeled samples and truly improve the disambiguation ability of $\mathcal{B}$, we summarize the transductive accuracy of $\mathcal{B}$ and $\mathcal{B}$-PLCP on real-world data sets and the results are shown in Table 7. It is obvious that $\mathcal{B}$-PLCP outperforms $\mathcal{B}$ in all cases according to the pairwise $t$-test with a significance level of 0.05, which validates that the disambiguation ability of $\mathcal{B}$ can be truly improved by PLCP. In other words, by integrating PLCP, classifiers can better identify and correct mislabeled samples, leading to outstanding performance.
Effectiveness of the Non-candidate Label Information
@@
where $\mathbf{W}$ and $\mathbf{b}$ are two model parameters and $\mathbf{L}\in\mathbb{R}^{n\times l}$ is the labeling confidence matrix. Denote this classifier $O$ and the original partner classifier $P$, the classification accuracy of PL-AGGD mutual-supervised with $O$ on each real-world data set is shown in the third row of Table 8. It is obvious that the performance of PL-AGGD coupled with $O$ is significantly inferior to that with $P$ in 87.5% cases, which validates the usefulness of the non-candidate label information. The accurate non-candidate label information provided by the partner classifier effectively complements the base classifier, leading to better performance.
Usefulness of the Blurring Mechanism
@@
It is also interesting to investigate the usefulness of the blurring mechanism in PLCP. By simply normalizing $\mathbf{P}_{1}$ to $\mathbf{O}_{1}$ and $\hat{\mathbf{P}}_{1}$ to $\hat{\mathbf{O}}_{1}$ (i.e., skipping Eq. (3) and Eq. (8)), the classification accuracy of PLCP without this mechanism is recorded in the second row of Table 8, where we can find that the performance of PL-AGGD-PLCP w/o Blur is significantly inferior to those with Blur in 87.5% cases. The blurring mechanism effectively tackles the overconfidence issue in PLL, leading to outstanding performance.
@@
Furthermore, we examine the impact of various values of $k$ on the performance of PLCP, as depicted in Fig. 1(d) in the Supplementary. It can be clearly observed that when $k$ is too small, the performance of PLCP deteriorates significantly. Additionally, the performance will be also exacerbated when $k\geq 0$, for in this case the predictions of the classifiers are enhanced rather than being blurred, amplifying small differences between two values. In this case, PLCP contributes little to disambiguation, resulting in inferior performance.
Influence of Kernel Extension
@@
We also conduct experiments to show the improvement of kernel extension used in partner classifier. Comparing the results in the first and second rows in Table 8, we can observe that the performance of PL-AGGD coupled with the classifier using kernel extension is superior to that without kernel extension on all the data sets, which validates the effectiveness of the kernel.
Sensitive Analysis

### v3

@@ 3.1 Compared Approaches
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News FG-NET(MAE3) FG-NET(MAE5)
PL-AGGD 0.063 $\pm$ 0.010 0.690 $\pm$ 0.020 0.451 $\pm$ 0.023 0.610 $\pm$ 0.012 0.521 $\pm$ 0.004 0.605 $\pm$ 0.002 0.418 $\pm$ 0.020 0.562 $\pm$ 0.020
@@
0.073 $\pm$ 0.011 $\bullet$ 0.698 $\pm$ 0.023 $\bullet$ 0.380 $\pm$ 0.013 $\bullet$ 0.542 $\pm$ 0.013 $\bullet$ 0.492 $\pm$ 0.003 $\bullet$ 0.463 $\pm$ 0.002 $\bullet$ 0.421 $\pm$ 0.020 $\bullet$ 0.560 $\pm$ 0.016 $\bullet$
@@
0.073 $\pm$ 0.006 $\bullet$ 0.721 $\pm$ 0.024 $\circ$ 0.471 $\pm$ 0.016 $\bullet$ 0.664 $\pm$ 0.012 $\bullet$ 0.521 $\pm$ 0.004 $\bullet$ 0.608 $\pm$ 0.003 $\bullet$ 0.422 $\pm$ 0.030 $\bullet$ 0.566 $\pm$ 0.020 $\bullet$
$O$ 0.071 $\pm$ 0.001 $\bullet$ 0.721 $\pm$ 0.004 $\circ$ 0.470 $\pm$ 0,020 $\bullet$ 0.663 $\pm$ 0.011 $\bullet$ 0.522 $\pm$ 0.003 $\bullet$ 0.605 $\pm$ 0.002 $\bullet$ 0.417 $\pm$ 0.022 $\bullet$ 0.576 $\pm$ 0.014 $\bullet$
$P$ 0.076 $\pm$ 0.009 0.717 $\pm$ 0.020 0.474 $\pm$ 0.015 0.668 $\pm$ 0.014 0.534 $\pm$ 0.003 0.611 $\pm$ 0.002 0.442 $\pm$ 0.018 0.582 $\pm$ 0.011
TABLE V: Ablation study of PLCP coupled with PICO on CIFAR-10 and CIFAR-100. $\bullet$/$\circ$ indicates whether PICO-PLCP is statistically superior/inferior to its degenerated version according to pairwise $t$-test at significance level of 0.05.
Partner Blur CIFAR-10 CIFAR-100
$q=0.1$ $q=0.3$ $q=0.5$ $q=0.01$ $q=0.05$ $q=0.1$
$P$
$\times$
94.60 $\pm$ 0.09 % $\bullet$ 94.20 $\pm$ 0.11 % $\bullet$ 93.63 $\pm$ 0.09 % $\bullet$ 73.71 $\pm$ 0.28 % $\bullet$ 73.17 $\pm$ 0.22 % $\bullet$ 69.91 $\pm$ 0.18 % $\bullet$
$O$ 94.44 $\pm$ 0.10 % $\bullet$ 94.37 $\pm$ 0.11 % $\bullet$ 93.59 $\pm$ 0.13 % $\bullet$ 73.43 $\pm$ 0.24 % $\bullet$ 72.98 $\pm$ 0.24 % $\bullet$ 69.77 $\pm$ 0.33 % $\bullet$
PICO-PLCP 94.80 $\pm$ 0.07 % 94.53 $\pm$ 0.10 % 93.67 $\pm$ 0.16 % 73.90 $\pm$ 0.20 % 73.51 $\pm$ 0.21 % 70.00 $\pm$ 0.35 %
3.2 Comparison with Stand-alone Methods
3.2.1 Experimental Settings
@@
We conduct experiments on six real-world partial label data sets collected from several domains and tasks, including FG-NET [14] for facial age estimation, Lost [18], Soccer Player [11] and Yahoo!News [12] for automatic face naming, MSRCv2 [25] for object classification and Mirflickr [26] for web image classification. For object classification task, image segmentations are taken as instances with objects appearing in the same image as candidate labels. For web image classification task, each example consists of images and a set of candidate tags extracted from the web page. The details of the data sets are summarized in Table I.
3.2.2 Performance on Real-World Data Set
It is noteworthy that the number of average label of FG-NET is quite large, which could cause quite low classification accuracy for all approaches. The common strategy is to evaluate the mean absolute error (MAE) between the predicted age and the ground-truth one. Specifically, we add another two sets of comparisons on FG-NET w.r.t. MAE3/MAE5, which means that test samples can be considered to be correctly classified if the difference between the predicted age and true age is no more than 3/5 years. Table II summaries the classification accuracy with standard deviation of each approach on real-world data sets, where we can observe that
@@ 3.2.2 Performance on Real-World Data Set
  • State-of-the-art (SOTA) approaches, such as PL-CL, PL-AGGD, LALO and SURE, can also be significantly improved by PLCP on all data sets. For instance, on FG-NET the performance of SURE can be improved by 45%, and on FG-NET(MAE3) the performance of LALO can be improved by 5%. Additionally, on Lost PL-AGGD can be improved by 4% and PL-CL can be improved by 5.5%. The partner classifier’s non-candidate label information effectively and significantly aids disambiguation, leading to outstanding performance of the PLCP framework.
@@
3.3 Comparison with Deep-learning Based Methods
3.3.1 Experimental Settings
In PLCP, we set $k=-1$ and $\mu=0.5$. All the hyper-parameters and experimental setups of $\mathcal{B}$ are set based on their original papers.
We conduct experiments on two benchmarks CIFAR-10 and CIFAR-100 [27], and following the settings in [23, 24], we generate false positive labels by flipping negative labels $\hat{y}\neq y$ for each sample with a probability $q=P(\hat{y}\in\mathcal{Y}|\hat{y}\neq y)$, where $y$ is the ground-truth label. In other words, the probability of a false positive is uniform across all $l-1$ negative labels. We combine the flipped ones with the ground-truth label to create the set of candidate labels. Specifically, $q$ is set to 0.1, 0.3 and 0.5 for CIFAR-10 and 0.01, 0.05 and 0.1 for CIFAR-100. Five independent runs are performed with the the average accuracy and standard deviation recorded (with different seeds).
3.3.2 Performance on CIFAR-10 and CIFAR-100
Table III presents the classification accuracy with standard deviation of each approach on CIFAR-10 and CIFAR-100, where we can find that
@@ 3.3.2 Performance on CIFAR-10 and CIFAR-100
3.4 Ablation Study
3.4.1 Effectiveness of the Non-candidate Label Information
@@
where $\mathbf{W}$ and $\mathbf{b}$ are two model parameters and $\mathbf{L}\in\mathbb{R}^{n\times l}$ is the labeling confidence matrix. Denote this classifier $O$ and the original partner classifier $P$, the classification accuracy of PL-AGGD mutual-supervised with $O$ on each real-world data set is shown in Table IV, and that of PICO on CIFAR data set is shown in Table V. It is obvious that the performances of PL-AGGD and PICO coupled with $O$ are significantly inferior to that with $P$ in 87.5% and all cases respectively, which validates the usefulness of the non-candidate label information. The accurate non-candidate label information provided by the partner classifier effectively complements the base classifier, leading to better performance.
Figure 4: Sensitivity of PLCP.
TABLE VI: Correction ratio of each method on real-world data sets.
Approaches Data set
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News FG-NET(MAE3) FG-NET(MAE5)
PL-CL 3.65 $\pm$ 0.02 % 21.29 $\pm$ 1.04 % 6.24 $\pm$ 0.48 % 7.96 $\pm$ 1.19 % 0.77 $\pm$ 0.08 % 4.19 $\pm$ 0.42 % 1.82 $\pm$ 0.19 % 2.46 $\pm$ 0.28 %
PL-AGGD 3.64 $\pm$ 0.04 % 20.00 $\pm$ 1.55 % 5.17 $\pm$ 0.49 % 9.31 $\pm$ 1.35 % 1.39 $\pm$ 0.21 % 1.99 $\pm$ 0.13 % 3.92 $\pm$ 0.84 % 4.81 $\pm$ 1.02 %
SURE 6.11 $\pm$ 0.02 % 6.67 $\pm$ 0.44 % 3.93 $\pm$ 0.73 % 11.39 $\pm$ 2.11 % 1.83 $\pm$ 0.26 % 7.45 $\pm$ 0.39 % 26.30 $\pm$ 2.48 % 34.90 $\pm$ 1.26 %
LALO 3.46 $\pm$ 0.01 % 13.77 $\pm$ 1.72 % 1.94 $\pm$ 0.35 % 7.43 $\pm$ 0.91 % 1.41 $\pm$ 0.22 % 3.48 $\pm$ 0.19 % 4.80 $\pm$ 1.26 % 5.85 $\pm$ 0.23 %
PL-SVM 7.49 $\pm$ 0.21 % 64.89 $\pm$ 2.51 % 31.28 $\pm$ 2.19 % 38.86 $\pm$ 2.47 % 28.10 $\pm$ 0.73 % 34.80 $\pm$ 0.88 % 38.98 $\pm$ 1.92 % 47.65 $\pm$ 2.59 %
PL-KNN 8.42 $\pm$ 0.17 % 56.93 $\pm$ 1.93 % 34.46 $\pm$ 1.83 % 49.81 $\pm$ 2.19 % 6.81 $\pm$ 0.52 % 43.47 $\pm$ 0.81 % 41.74 $\pm$ 2.01 % 39.22 $\pm$ 1.82 %
TABLE VII: Transductive accuracy of each compared approach on the real-world data sets. For any compared approach $\mathcal{B}$, $\bullet$/$\circ$ indicates whether $\mathcal{B}$-PLCP is statistically superior/inferior to $\mathcal{B}$ according to pairwise $t$-test at significance level of 0.05.
Approaches Data set
FG-NET Lost MSRCv2 Mirflickr Soccer Player Yahoo!News FG-NET(MAE3) FG-NET(MAE5)
PL-CL 0.159 $\pm$ 0.016 0.832 $\pm$ 0.019 0.585 $\pm$ 0.012 0.697 $\pm$ 0.019 0.715 $\pm$ 0.001 0.827 $\pm$ 0.003 0.600 $\pm$ 0.029 0.737 $\pm$ 0.018
PL-CL-PLCP 0.181 $\pm$ 0.012 $\bullet$ 0.854 $\pm$ 0.010 $\bullet$ 0.638 $\pm$ 0.008 $\bullet$ 0.704 $\pm$ 0.021 $\bullet$ 0.722 $\pm$ 0.004 $\bullet$ 0.831 $\pm$ 0.003 $\bullet$ 0.618 $\pm$ 0.023 $\bullet$ 0.747 $\pm$ 0.018 $\bullet$
PL-AGGD 0.141 $\pm$ 0.012 0.793 $\pm$ 0.020 0.557 $\pm$ 0.015 0.695 $\pm$ 0.015 0.669 $\pm$ 0.003 0.808 $\pm$ 0.005 0.571 $\pm$ 0.027 0.709 $\pm$ 0.025
PL-AGGD-PLCP 0.167 $\pm$ 0.015 $\bullet$ 0.829 $\pm$ 0.016 $\bullet$ 0.638 $\pm$ 0.011 $\bullet$ 0.715 $\pm$ 0.015 $\bullet$ 0.714 $\pm$ 0.004 $\bullet$ 0.831 $\pm$ 0.004 $\bullet$ 0.600 $\pm$ 0.022 $\bullet$ 0.738 $\pm$ 0.016 $\bullet$
SURE 0.158 $\pm$ 0.012 0.796 $\pm$ 0.026 0.603 $\pm$ 0.016 0.650 $\pm$ 0.024 0.700 $\pm$ 0.003 0.798 $\pm$ 0.005 0.590 $\pm$ 0.019 0.727 $\pm$ 0.020
SURE-PLCP 0.170 $\pm$ 0.011 $\bullet$ 0.836 $\pm$ 0.021 $\bullet$ 0.622 $\pm$ 0.015 $\bullet$ 0.697 $\pm$ 0.017 $\bullet$ 0.706 $\pm$ 0.005 $\bullet$ 0.827 $\pm$ 0.006 $\bullet$ 0.605 $\pm$ 0.026 $\bullet$ 0.739 $\pm$ 0.018 $\bullet$
LALO 0.153 $\pm$ 0.017 0.818 $\pm$ 0.019 0.548 $\pm$ 0.009 0.681 $\pm$ 0.013 0.688 $\pm$ 0.004 0.822 $\pm$ 0.004 0.593 $\pm$ 0.025 0.730 $\pm$ 0.015
LALO-PLCP 0.168 $\pm$ 0.018 $\bullet$ 0.831 $\pm$ 0.019 $\bullet$ 0.620 $\pm$ 0.009 $\bullet$ 0.694 $\pm$ 0.019 $\bullet$ 0.706 $\pm$ 0.004 $\bullet$ 0.827 $\pm$ 0.004 $\bullet$ 0.604 $\pm$ 0.025 $\bullet$ 0.741 $\pm$ 0.019 $\bullet$
PL-SVM 0.176 $\pm$ 0.015 0.609 $\pm$ 0.055 0.570 $\pm$ 0.040 0.581 $\pm$ 0.022 0.660 $\pm$ 0.008 0.691 $\pm$ 0.005 0.566 $\pm$ 0.025 0.706 $\pm$ 0.024
PL-SVM-PLCP 0.188 $\pm$ 0.008 $\bullet$ 0.792 $\pm$ 0.036 $\bullet$ 0.639 $\pm$ 0.027 $\bullet$ 0.630 $\pm$ 0.028 $\bullet$ 0.705 $\pm$ 0.003 $\bullet$ 0.823 $\pm$ 0.004 $\bullet$ 0.580 $\pm$ 0.027 $\bullet$ 0.721 $\pm$ 0.019 $\bullet$
PL-KNN 0.041 $\pm$ 0.007 0.337 $\pm$ 0.030 0.415 $\pm$ 0.014 0.466 $\pm$ 0.013 0.493 $\pm$ 0.004 0.403 $\pm$ 0.010 0.285 $\pm$ 0.017 0.438 $\pm$ 0.015
PL-KNN-PLCP 0.168 $\pm$ 0.012 $\bullet$ 0.781 $\pm$ 0.033 $\bullet$ 0.639 $\pm$ 0.021 $\bullet$ 0.628 $\pm$ 0.015 $\bullet$ 0.699 $\pm$ 0.006 $\bullet$ 0.793 $\pm$ 0.004 $\bullet$ 0.579 $\pm$ 0.025 $\bullet$ 0.722 $\pm$ 0.018 $\bullet$
3.4.2 Usefulness of the Blurring Mechanism
@@
It is also interesting to investigate the usefulness of the blurring mechanism in PLCP. By simply normalizing $\mathbf{P}_{1}$ to $\mathbf{O}_{1}$ and $\hat{\mathbf{P}}_{1}$ to $\hat{\mathbf{O}}_{1}$ (i.e., skipping Eq. (3) and Eq. (8)), the classification accuracy of PLCP without this mechanism is recorded in Table IV-V, where we can find that the performance of PL-AGGD-PLCP and PICO-PLCP w/o Blur is significantly inferior to those with Blur in 87.5% and all cases respectively. The blurring mechanism effectively tackles the overconfidence issue in PLL, leading to outstanding performance.
@@
Furthermore, we examine the impact of various values of $k$ on the performance of PLCP, as depicted in Fig. 4(d). It can be clearly observed that when $k$ is too small, the performance of PLCP deteriorates significantly. Additionally, the performance will be also exacerbated when $k\geq 0$ (strictly $k>\ln{2}$), for in this case the predictions of the classifiers may be enhanced rather than being blurred, amplifying small differences between two values. As discussed in Sec. 4.3, PLCP contributes little to appeal and disambiguation in this situation, resulting in inferior performance.
3.4.3 Influence of Kernel Extension
@@
We also conduct experiments to show the improvement of kernel extension used in partner classifier. Comparing the results in the first and second rows in Table IV, we can observe that the performance of PL-AGGD coupled with the classifier using kernel extension is superior to that without kernel extension on all the data sets, which validates the effectiveness of the kernel.
3.4.4 Sensitive Analysis